C= 5(f-32)divided by 9 for f
Answer by Alan3354(69443)

C = (5/9)*(F-32)
(9/5)*C = F - 32
F = (9/5)*C + 32
----------------
F = (9*C + 160)*2/10 more useful
eg, to convert from 37 C to F:
37*9 = 333
Add 160 = 493
x2 = 986
move the decimal --> 98.6 F commonly used as normal body temperature, tho there's at least 1 deg F variation in that.